# Gaze3P: Gaze-Based Prediction of User-Perceived Privacy

### v1

@@ 5.1. Stimuli-based Perceived Privacy (SPP)
When data minimisation constraints are not imposed, we assess model performance with the inclusion of stimulus data and compare it to PrivacEye (Steil et al., 2019). Results in Table 1-gaze+stimuli show that the stimulus features significantly improve the performance.
@@ 5.2. User-based Perceived Privacy (UPP)
UPP applications.  User-specific insights can, for example, be used for privacy nudges (Ioannou et al., 2021) (to encourage users to make privacy-conscious decisions according to their expertise or profiles, commonly used in social media platforms to prevent oversharing, e.g. Facebook’s contextual privacy warnings (Meta, 2024) and Chrome browser-based security warnings (Google, 2025)), adaptive two-factor-authentication (2FA) (Djeki et al., 2024) (which requires extra authentication only in suspicious situations), privacy labels and transparency notices (to help users understand and control their privacy choices according to their profiles, e.g. Apple’s App Store privacy labels (Apple, 2025)), and cohort-based recommendations (for group-based personalization, e.g. Google’s Federated Learning of Cohorts (FLoC) (Google, 2021) for ad targeting).
@@
The difference in gaze behaviour between experts and non-experts arises due to cognitive and perceptual adaptations that develop with experience and training (Brams et al., 2019; Liu et al., 2009; Law et al., 2004). Our results show that privacy experts exhibit more focused attention on privacy-relevant information and quicker identification of potential privacy risks compared to non-experts. In other words, privacy experts have rapid gaze shifts, suggesting automatic heuristic-based decision-making, while non-experts have longer dwell time, indicating uncertainty and high cognitive load in privacy assessment (tested by a Mann-Whitney U-test with $p\text{-}value=$0.03$<$0.05$$), leading to distinguishing features that ML models can identify with high accuracy (c.f. Table 3).
@@ 6.1. Privacy-Utility-Cognition Trade-Off
Since users do not select privacy levels purely based on objective risks or technical parameters, but rather through subjective interpretations shaped by cognitive biases (cf. Appendix B), we now want to discuss the typical choices of $g$ in cognitive science and how they affect the accuracy of a model obfuscated with $\varepsilon$-DP noise if $\varepsilon=f(l)$.
@@
It is important to note that users employ diverse rationales and cognitive strategies when forming their privacy perceptions, and these strategies are shaped by multiple interacting factors, such as the stimuli, background knowledge, and demographic (cf. Section 5). As a result, there is no universally correct mapping function, as these differences reflect valid subjective interpretations rooted in cognitive and contextual variability. Rather than assuming a single ground truth, such mappings should be empirically learned from user behaviour and validated through data-driven analysis. Hence, later in this section, we illustrate this point by providing a practical example based on our dataset.
@@ 6.1.1. Linear Mapping
This mapping assumes equal intervals between privacy levels. Linear mappings are widely used in statistics, cognitive sciences, and machine learning for their mathematical simplicity, speed, and interoperability, and serve as good approximation functions. Nonetheless, they may potentially lead to inaccurate interpretations (Bürkner and Vuorre, 2019). This is because, although privacy levels are numerically treated as equidistant, the psychological or cognitive perception of the distance between those points may not be equal (Waldman, 2020; Kezer et al., 2022). For example, the difference between ’very safe’ and ’safe’ may be much smaller in the user’s mind than the difference between ’neutral’ and ’moderately private’. Similarly, the midpoint (’neutral’) may be seen not as a numerical centre but rather as a safe choice. Hence, we now longer have a linear relation.
@@ 6.2. Empirical Evaluation
We now show how our results from Section 5 could be used in combination with mappings to epsilon values proposed in Section 6.1. To illustrate the applicability of our approach, we adopt a representative example from the privacy level perception task. This example serves to show how our method can be operationalized in practice. Importantly, the proposed framework is not limited to this specific task; it can be extended to other tasks by incorporating stimuli-or user-based information, thereby enhancing its generalizability across various privacy-sensitive applications.
We apply the state-of-the-art personalised differential privacy (PDP) mechanisms from (Jorgensen et al., 2015) for privacy budgets $\varepsilon$ determined with our mapping functions from Section 6.1. Recall that (Jorgensen et al., 2015) is originally used with randomly chosen $\varepsilon$. For our comparison, we use different common tasks from data analysis and machine learning tasks, which we want to describe briefly next.
@@
Our evaluation uses the same benchmarks as Jorgensen et al. (Jorgensen et al., 2015) to evaluate PDP mechanisms for search analytics, e.g. count query for the number of documents that participants searched for, median and minimum queries for attention allocation (number of fixations) per stimulus. Once a $\varepsilon$-value is determined, either randomly by (Jorgensen et al., 2015) or by applying our method on the user’s gaze data, a trusted data analyst receives the data and the $\varepsilon$ values, then adds suitable noise to query results and publishes the aggregate statistics.
In addition to these classical data analysis benchmarks, we also evaluate PDP mechanisms for machine learning tasks, namely, search intention prediction, using the default models and parameters of (Jorgensen et al., 2015; Boenisch et al., 2023, 2022) (cf. Section 3).
@@ Results
As shown in Table 4, PDP approaches significantly improve over the static DP approaches (i.e. worst-case $\varepsilon$, the common approach in DP) and are significantly closer to the plain approaches (i.e. without noise addition). In addition, our proposed mapping functions yield better utility than the random benchmark (i.e. the random $\varepsilon$ values generated by the existing benchmarks).
@@
  • The sequential mapping demonstrates superior performance for attributes with deliberative privacy sensitivity ($52<range(l)<$ 5), such as political opinions, where participants’ gaze shifts between the privacy levels before selection, suggesting an initial classification of the attribute as private, followed by a secondary assessment of its degree of privacy.
@@ 7. Discussion
Our dataset showed that eye gaze reliably reflects user privacy perceptions across privacy-related tasks. It further revealed user- and stimulus-based insights that could be used across a variety of applications.
@@
By modelling perceived privacy alongside formal guarantees, we improved data utility while (cognitively) aligning with individual privacy expectations. Although a single universal mapping between perception and privacy budgets is unlikely to exist due to inter-individual variability, we propose a set of adaptable mapping functions that can be selected contextually based on application needs and behavioural insights.
@@
Our proposed datasets include further information that we encourage the community to develop upon. This can include (i) additional tasks, such as multi-attribute and multi-user correlation analyses (i.e. each group of stimuli was shown to various participants, with each stimulus containing several attributes), (ii) more advanced models to capture more patterns and enhance our baseline performance, and (iii) other DP and privacy-preserving protocols (i.e. adding noise to different privacy units, the fundamental entity whose privacy is protected, user-, label-, feature- or pixel-level DP) that would benefit from the user (cognitive) privacy perceptions without explicit interaction.
@@ Acknowledgements.
We thank Manpa Barman and Yanhong Xu for their assistance in recruiting participants. M. Elfares was funded by the Ministry of Science, Research and the Arts Baden-Württemberg in the Artificial Intelligence Software Academy (AISA). P. Reisert and R. Küsters were supported by the German Federal Ministry of Education and Research under Grant Agreement No. 16KIS1441 (CRYPTECS project) and the German Research Foundation under Grant No. 548713845.
@@ A.4. Data Collection
To accurately track participants’ gaze and record eye movement data during the experiment, we first conduct a 13-point calibration to ensure the eye tracker is accurately mapping gaze coordinates to the screen, followed by a validation check to confirm that gaze accuracy is within acceptable error limits (e.g., ¡0.5°). We further recalibrate if drift or errors occur during the session. The stimuli are then presented and data streams are recorded.
@@ A.5. Data Processing
We structured the processed data into a usable dataset for analysis, by organizing data into rows and columns ( .csv format), including information such as participant demographics, stimuli details, and task condition. Finally, we formatted the data for statistical or machine learning tools (e.g., Python) in the following structure:
@@
Appendix B Cognitive Theories and Privacy

### v2

@@ 5.1. Stimuli-based Perceived Privacy (SPP)
We assess model performance with the inclusion of stimulus data and compare it to PrivacEye (Steil et al., 2019). Results in Table 1-gaze+stimuli show that the stimulus features significantly improve the performance. Nonetheless, models trained solely on gaze data still yield meaningful results. This demonstrates that gaze behaviour alone carries informative signals related to perceived privacy, supporting its utility in scenarios where access to full visual content may be restricted or deliberately excluded for data minimization and privacy-preserving purposes.
Bottom-up and top-down attention
To minimise the confounding influence of bottom-up (saliency-driven) gaze, we attempted to control the visual composition of our stimuli by selecting the VISPR images that mostly contain a single dominant attribute. This design choice was intended to reduce the likelihood that gaze behaviour would be driven by reflexive visual attention to salient features, such as bright colours, known to automatically attract attention regardless of the observer’s cognitive or emotional state. Such bottom-up attention is inherently stimulus-driven and unrelated to higher-order constructs like perceived privacy (i.e. top-down attention). However, this assumption may not hold in all scenarios, especially when stimuli are real-world images. To address this, we check if residual gaze patterns — i.e., the part of gaze behaviour not explained by bottom-up saliency — are more predictive of perceived privacy than raw gaze data. We use a saliency prediction model (DeepGaze (Kümmerer et al., 2022)) to generate a saliency map for each stimulus image. This represents the expected gaze distribution if attention were purely bottom-up. Then we compute the residual gaze maps by subtracting the (normalised) saliency map from the original gaze map. We input both the original gaze maps as well as the residual gaze maps to the models, as shown in Table 1. Results are significantly lower for residual gaze maps, which may indicate that privacy-relevant content often coincides with visually salient regions. Additionally, the lack of variance in saliency features reduces the need for correction, leading to differences in performance between the two models, since the stimuli used in the study were relatively homogeneous in terms of saliency, mostly with one dominant object.
@@ 5.2. User-based Perceived Privacy (UPP)
UPP applications.  User-specific insights can, for example, be used for privacy nudges (Ioannou et al., 2021) (to encourage users to make privacy-conscious decisions according to their expertise or profiles, commonly used in social media platforms to prevent oversharing, e.g. Facebook’s contextual privacy warnings (Meta, 2024) and Chrome browser-based security warnings (Google, 2025)), privacy labels and transparency notices (to help users understand and control their privacy choices according to their profiles, e.g. Apple’s App Store privacy labels (Apple, 2025)), and cohort-based recommendations (for group-based personalization, e.g. Google’s Federated Learning of Cohorts (FLoC) (Google, 2021) for ad targeting) (C.f. Appendix C).
@@
The difference in gaze behaviour between experts and non-experts arises due to cognitive and perceptual adaptations that develop with experience and training (Brams et al., 2019; Liu et al., 2009; Law et al., 2004). Our results show that privacy experts exhibit more focused attention on privacy-relevant information and quicker identification of potential privacy risks compared to non-experts. In other words, privacy experts have rapid gaze shifts, suggesting automatic heuristic-based decision-making, while non-experts have longer dwell time, indicating uncertainty and high cognitive load in privacy assessment ($H_{0}$: ’The gaze duration distributions of the two groups (experts and non-experts) are equal’ tested by a Mann-Whitney U-test with $p\text{-}value=$0.03$<$0.05$\$ and effect size $r=0.3$), leading to distinguishing features that ML models can identify with high accuracy (c.f. Table 3).
@@ 6.1. Privacy-Utility-Cognition Trade-Off
Since users do not select privacy levels purely based on objective risks or technical parameters, but rather through subjective interpretations shaped by cognitive biases (cf. Appendix D), we now want to discuss the typical choices of $g$ in cognitive science and how they affect the accuracy of a model obfuscated with $\varepsilon$-DP noise if $\varepsilon=f(l)$.
@@
It is important to note that users employ diverse rationales and cognitive strategies when forming their privacy perceptions, and these strategies are shaped by multiple interacting factors, such as the stimuli, background knowledge, and demographics (cf. Section 5). As a result, there is no universally correct mapping function, as these differences reflect valid subjective interpretations rooted in cognitive and contextual variability. Rather than assuming a single ground truth, such mappings should be empirically learned from user behaviour and validated through data-driven analysis. Hence, later in this section, we illustrate this point by providing a practical example based on our dataset.
@@ 6.1.1. Linear Mapping
This mapping assumes equal intervals between privacy levels. Linear mappings are widely used in statistics, cognitive sciences, and machine learning for their mathematical simplicity, speed, and interoperability, and serve as good approximation functions. Nonetheless, they may potentially lead to inaccurate interpretations (Bürkner and Vuorre, 2019). This is because, although privacy levels are numerically treated as equidistant, the psychological or cognitive perception of the distance between those points may not be equal (Waldman, 2020; Kezer et al., 2022). For example, the difference between ’very safe’ and ’safe’ may be much smaller in the user’s mind than the difference between ’neutral’ and ’moderately private’. Similarly, the midpoint (’neutral’) may be seen not as a numerical centre but rather as a safe choice. Hence, we no longer have a linear relation.
@@ 6.2. Empirical Evaluation
We now show how our results from Section 5 could be used in combination with mappings to epsilon values proposed in Section 6.1. To illustrate the applicability of our approach, we adopt a representative example from the privacy level perception task. This example serves to show how our method can be operationalised in practice. Importantly, the proposed framework is not limited to this specific task; it can be extended to other tasks by incorporating stimuli-or user-based information, thereby enhancing its generalizability across various privacy-sensitive applications.
We apply the state-of-the-art personalised differential privacy (PDP) mechanisms for privacy budgets $\varepsilon$ determined with our mapping functions from Section 6.1. We employ the existing PDP methods without modification to their core mechanisms. The only alteration involves replacing the randomly generated privacy-level inputs ($\varepsilon$) commonly used in prior work with user-specified privacy preferences inferred from gaze. This allows for a more realistic and user-informed evaluation of PDP behaviour while preserving the original algorithmic structure. For our comparison, we use different common tasks from data analysis and machine learning tasks, which we want to describe briefly next.
@@
Our evaluation uses the same benchmarks as Jorgensen et al. (Jorgensen et al., 2015) to evaluate PDP mechanisms for search analytics, e.g. count query for the number of documents that participants searched for, median and minimum queries for attention allocation (number of fixations) per stimulus. Once a $\varepsilon$-value is determined, either randomly by (Jorgensen et al., 2015) or by applying our method on the user’s gaze data, a trusted data analyst receives the data and the $\varepsilon$ values, then adds suitable noise to query results and publishes the aggregate statistics. In addition to these classical data analysis benchmarks, we also evaluate PDP mechanisms for machine learning tasks, namely, search intention prediction, using the default models and parameters of (Jorgensen et al., 2015; Boenisch et al., 2023, 2022) (cf. Section 3).
@@ Results
As shown in Table 4, PDP approaches significantly improve over the static DP approaches (i.e. the worst-case privacy guarantee by uniformly applying the maximum privacy level required across all users, without adapting to individual preferences or contexts to address the most stringent privacy demands, the common approach in DP) and are significantly closer to the plain approaches (i.e. without noise addition). In addition, our proposed mapping functions yield better utility than the random benchmark (i.e. the random $\varepsilon$ values generated by the existing benchmarks).
@@
  • The sequential mapping demonstrates superior performance for attributes with deliberative privacy sensitivity ($2<range(l)<5$), such as political opinions, where participants’ gaze shifts between the privacy levels before selection, suggesting an initial classification of the attribute as private, followed by a secondary assessment of its degree of privacy.
@@
User expectation alignment
To assess whether the predictions and protections provided by Gaze3P align with user expectations, we conducted a follow-up user validation study with a subset of $N=20$ participants. While our quantitative evaluations above demonstrate the technical validity of Gaze3P (e.g., accuracy), the user study examines the users’ subjective acceptance of Gaze3P.
For the validation study, each participant was shown the same set of images they had rated previously during the initial dataset collection (the search task), along with their user-specified privacy levels. For each image, we generated a reconstructed version using the standard reconstruction attack pipeline from MLDoctor (Liu et al., 2022b) to attack the PDP learning models, simulating an adversarial attempt to recover visual data from differentially private representations. These reconstructions, shown in Fig. 8, reflect what an external observer might infer about the original image when Gaze3P’s predicted privacy levels are used in PDP.
Participants were asked three questions measuring (i) utility, (ii) privacy, and (iii) cognition: (i) if the model’s search target prediction was correct. Responses were binary (’yes’ or ’no’). (ii) whether the reconstructed image matched their privacy expectations, given the privacy level they previously selected. Responses were collected on a 5-point Likert scale, ranging from ’Not at all’ to ’Perfectly aligned’, and (iii) if they prefer a certain mapping function, given this privacy-utility tradeoff. Responses were again collected on a 5-point Likert scale per mapping function. We repeated this procedure for each stimulus, leading to each participant being shown $250$ samples (the same $50$ search stimuli that were previously shown during data collection times our $5$ benchmark models).
Results show that the subsample shown to participants was representative, and utility evaluation was close to the numbers reported in Table 4 with 0.5 (i.e. 50%), 0.6, 0.5, 0.7, and 0.5 correct predictions on average for the learning models in the same order. In addition, with an average rating of $3.8/5$, the participants’ privacy expectations were met. Finally, overall, participants preferred the sequential mapping in 68% of the cases. They preferred the exponential mapping in 85% of the private images (rated with $l\geq 6$) and the sequential mapping in 52% of the images rated as ’neutral’, which supports our quantitative results. However, the sigmoid mapping was only selected in 7% of the cases, showing no clear pattern.
Figure 8. Random samples of the reconstructed stimuli (first row $l=7$, second row $l=1$.
@@ 7. Discussion
The introduction of Gaze3P marks a pioneering step in privacy perception, advancing beyond traditional questionnaires by leveraging implicit, dynamic signals derived from gaze behaviour, underscoring that users’ visual attention can implicitly reveal privacy sensitivities—offering a real-time, non-intrusive window into subjective privacy evaluation. More concretely, our dataset showed that eye gaze reliably reflects user perceptions across privacy-related tasks (RQ1). Our ML approach further revealed user- and stimulus-based insights that could be used across a variety of applications (RQ2): We demonstrated the predictive power of gaze for privacy perception, the behavioural and contextual influences of gaze (e.g. fixation patterns, visual attention, demographic context, and expertise) in shaping privacy perceptions, the potential of gaze-based inference and enhancement of privacy-related attribute recognition and user profiling.
@@
By modelling perceived privacy alongside formal guarantees, we improved data utility while (cognitively) aligning with individual privacy expectations (RQ3). Although a single universal mapping between perception and privacy budgets is unlikely to exist due to inter-individual variability, we propose a set of adaptable mapping functions that can be selected contextually based on application needs and behavioural insights.
7.1. Limitations
We acknowledge an age-related bias in our data, as the participant pool was not age-controlled. Despite a publicly announced call, recruitment likely skewed towards university students and staff due to convenience and accessibility. We also collected supplementary standard demographics, including education level, occupation, and field of study. However, our analysis revealed limited variability across these dimensions, as the majority of participants were students enrolled in STEM (Science, Technology, Engineering, Mathematics) disciplines, with only a few exceptions (4 participants). Consequently, while the current dataset may not support strong generalization across diverse demographic strata, this metadata, as well as our data collection software, is included in (Author, s). We hope that this allows others to extend analyses or augmentations in future work, particularly with a more demographically diverse sample.
Moreover, participants were explicitly instructed to treat the images as if they were their own (e.g., from a personal phone gallery), in line with Sattar et al.  (Sattar et al., 2020). They were also informed that the study focused on privacy, in accordance with the ethical guidelines and prior research practices (Orekondy et al., 2017; Steil et al., 2019). While we acknowledge the limitations of this setup and the possible biases it may induce, such constraints are typical in early-stage studies within emerging research domains. Nonetheless, this work serves as a foundational step, enabling future research to adopt more comprehensive methodologies and engage more diverse participant groups to improve generalizability and depth of analysis.
Similarly, another limitation of the present study is the introduction of contextual bias through priming participants on the topic of privacy. This framing was necessary to align participants with the experimental objectives; however, it likely influenced cognitive processing and visual attention during the tasks. As a result, our predictions are not solely inferred from gaze behaviour in a neutral context but rather from gaze patterns shaped by a privacy-salient environment. This contextualization restricts the generalizability of the findings, as gaze allocation strategies may differ when privacy is not explicitly emphasised. Consequently, the reported predictive performance should be interpreted as applying to privacy-aware scenarios, not to all interaction contexts. Future research should address this by employing between-subject designs with primed and non-primed conditions or by modelling contextual factors explicitly, to disentangle intrinsic gaze-based indicators of privacy perception from those induced by experimental framing.
7.2. Future Work
In this paper, we showed that perceived privacy preferences can be inferred from gaze behaviour alone. While not the main focus of the present analysis, we also observed that contextual data, such as demographics, have the potential to enrich predictions. Future work can build on this foundation by incorporating such auxiliary information to enhance both accuracy and personalisation. Similarly, additional metadata on individual user traits, such as trust propensity, risk perception, and prior exposure to privacy threats, can be collected to enable a deeper understanding of how such latent factors implicitly influence privacy-related behaviours (Guerra et al., 2023b; Prange et al., 2024; Elbitar et al., 2021).
@@ 7.2. Future Work
Additionally, our proposed dataset includes rich information that we encourage the community to develop upon. This can include (i) additional tasks, such as multi-attribute and multi-user correlation analyses (i.e. each group of stimuli was shown to various participants, with each stimulus containing several attributes), (ii) tailored models to capture more patterns and enhance our baseline performance, (iii) other DP and privacy-preserving protocols (i.e. adding noise to different privacy units, the fundamental entity whose privacy is protected, user-, label-, feature- or pixel-level DP) that would benefit from the user (cognitive) privacy perceptions without explicit interaction, and (iv) fine-grained eye-tracking analyses to gain a deeper understanding of gaze behaviour in privacy-sensitive contexts. While the present study demonstrates that learned gaze features can serve as effective predictors of perceived privacy, more granular analyses - e.g. gaze entropy, micro-saccade dynamics, and scanpath structure - may reveal subtle cognitive and affective processes underlying privacy perception and could help disentangle the interplay between bottom-up and top-down visual attention in private settings, improving the interpretability of gaze-based privacy models.
@@ Acknowledgements.
We thank Manpa Barman and Yanhong Xu for their assistance in recruiting participants. M. Elfares was funded by the Ministry of Science, Research and the Arts Baden-Württemberg in the Artificial Intelligence Software Academy (AISA).
@@ A.4. Data Collection
To accurately track participants’ gaze and record eye movement data during the experiment, we first conduct a 13-point calibration to ensure the eye tracker is accurately mapping gaze coordinates to the screen, followed by a validation check to confirm that gaze accuracy is within acceptable error limits (e.g., ¡0.5°). We further recalibrate if drift or errors occur during the session. The stimuli are then presented, and data streams are recorded.
Essential methodological safeguards were implemented to maintain participant engagement and ensure high data quality throughout the experiment. These safeguards served to minimise bias, reduce participant fatigue, and confirm that responses reflected genuine attention and comprehension.
Randomised attention checks were embedded at multiple points during the study. These checks consisted of stimuli with clearly identifiable attributes—some explicitly private (e.g., credit cards) and others non-private (e.g., randomly generated colour patches). Participants were asked to classify or respond to these items. Correct responses indicated attentiveness, while incorrect answers were flagged for potential disengagement or misunderstanding.
To mitigate cognitive fatigue, structured breaks were introduced at predefined intervals (i.e. between tasks), allowing participants to rest and maintain focus. Participants were also allowed to stop the experiment at any point, if needed.
To address order and learning effects, both the sequence of experimental tasks and the presentation order of stimuli were fully randomised for each participant. This ensured that performance patterns could not be attributed to predictable task progression, practice effects, or fatigue tied to task order.
All these procedures align with established best practices in behavioural and user research (Carter and Luke, 2020; Hessels et al., 2025; Hooge et al., 2025; Nyström et al., 2025; Niehorster et al., 2025), ensuring internal validity while safeguarding the participant experience.
@@ A.5. Data Processing
We structured the processed data into a usable dataset for analysis by organizing data into rows and columns ( .csv format), including information such as participant demographics, stimuli details, and task condition. Finally, we formatted the data for statistical or machine learning tools (e.g., Python) in the following structure:
@@
Appendix B ML Implementation Details
Given a dataset of inputs (e.g. gaze data) annotated with labels (e.g. user-selected privacy levels), we define the training set as:
$\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$
where: $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the gaze data of the $i^{\text{th}}$ sample (or a feature vector extracted from the stimuli and the gaze when both are inputs, as defined by each task), $y_{i}\in\mathcal{Y}=\{1,2,\dots,K\}$ is the corresponding label (e.g. user-provided privacy level or attribute name or privacy expertise, as defined by each task).
The training phase:
In a supervised learning setting like classification, the training phase uses labelled data—pairs of input features ($x_{i}$) and their corresponding ground truth labels ($y_{i}$). During training, the model learns a mapping from inputs to labels by minimising a loss function that measures the discrepancy between predicted and true labels.
The goal is to learn a function $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathcal{Y}$, parameterized by $\theta$, that minimizes the classification loss:
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{\theta}(\mathbf{x}_{i}),y_{i})$
where $\ell(\cdot,\cdot)$ is the loss function.
The inference phase:
During the inference (testing) phase, the model receives only input features from new, unseen samples and generates predicted labels based on the patterns it learned during training. No true labels are available during inference; predictions are made autonomously using the trained model parameters.
More technically, a new unseen gaze sample $\mathbf{x}^{\ast}\in\mathbb{R}^{d}$ is processed through the trained model to produce a predicted label:
$\hat{y}=f_{\theta^{\ast}}(\mathbf{x}^{\ast})$
We then compare the predictions $\hat{y}_{i}$ with the true labels $y_{i}$ only to quantify the models’ performance.
Train-test split:
We employ cross-validation to optimise generalisation performance and prevent overfitting. In cross-validation, the data is split into $k$ folds; the model is trained on k1k-1 folds and tested on the remaining one, repeating this process $k$ times so that each fold serves as a test set once. The final performance is averaged across all folds, providing a more robust and unbiased estimate than a single train-test split. Here, we used an 80-20 train-test split.
ML models:
In our implementation, we mostly employed classical machine learning algorithms available within the Scikit-learn framework, using their default configurations. Decision Trees were implemented using ’DecisionTreeClassifier’. Support Vector Machines (SVMs) were used via ’SVC’ with an ’RBF kernel’. Logistic Regression was applied using ’LogisticRegression’ with L2 regularisation. For Random Forests, we utilised ’RandomForestClassifier’, an ensemble of decision trees trained on bootstrapped subsets of data with feature bagging. K-Nearest Neighbors (KNN) was implemented using ’KNeighborsClassifier’, which assigns class labels based on the majority vote among the k most similar training samples in feature space. Further information and default hyperparameters can be found in the Scikit-Learn documentation: scikit-learn.org/stable/supervised_learning.html. Finally, for other advanced models, we used the original implementations provided by the referenced papers.
Appendix C Potential Real-World Applications
In this section, we discuss the potential SPP and UPP applications in more details:
SPP applications.  Stimuli-based perceived privacy results can be used in applications such as:
  • Setting hand-picked parameters in privacy-preserving protocols: Similar to our DP application in Section 6, federated learning (FL) (McMahan et al., 2017; Kairouz et al., 2021) can also benefit from our approach. For example, the model update perturbation step involves adding random noise to the model updates (usually gradients or weights) before sending them to the central server. This includes noise scale $\sigma$ and a clipping norm $c$ parameters. These are typically hand-picked through empirical tuning or heuristics to achieve a privacy-utility tradeoff. Hence, Gaze3P can be used to implicitly set these parameters and personalise them to the users’ expectations. The same applies to K-anonymity (Samarati and Sweeney, 1998), L-diversity (Machanavajjhala et al., 2007), T-closeness (Li et al., 2006), to set the $k$, $L$, and $T$ values.
  • Privacy auctions: Privacy auctions (Ghosh and Roth, 2011; Zhang et al., 2020) are mechanisms where users ”sell” their private data or privacy loss in exchange for compensation. These auctions aim to determine how much privacy loss users are willing to tolerate, and at what cost, allowing systems to personalise privacy levels across individuals based on their preferences. Users specify a subjective cost or price that they associate with a unit of privacy loss, typically through surveys. Hence, Gaze3P offers a more robust and implicit feedback with the need for users to understand and evaluate their data.
  • Synthetic data generation: Synthetic data generation (Liu et al., 2022a) refers to the creation of artificial data that mimics the statistical properties of real user data, often used to preserve privacy while enabling data analysis or model training. Here, again Gaze3P can be used to match the user privacy expectations with respect to the data, e.g. data fidelity preferences (acceptable levels of distortion or allowed deviation from real data distributions).
  • Access control models: Access control models regulate who can access what data under which conditions. Two key types are: (i) Attribute-Based Access Control (ABAC) (Hu et al., 2015), which dynamically grants or denies access based on user, resource, and environmental attributes (e.g., role, location, time). The stimuli-specific insights of Gaze3P, can be used to automatically infer these attributes (e.g. the private attribute recognition task) and set the corresponding privacy parameters (e.g. access) without explicitly defining and listing all possible roles (e.g. the privacy expertise or the user profiling tasks). (ii) Human-in-the-loop privacy controls, which empower users to manage privacy interactively in real-time systems. In this case, Gaze3Pcan be integrated with, for example, Instagram’s ’Restrict’ feature (Parmelee and Roman, 2020) or e-mail spam filters to allow users to flag content without explicitly selecting buttons and going through lists.
UPP applications.  User-specific insights can also be integrated into several real-world applications such as:
  • Privacy nudges: Privacy nudges (Ioannou et al., 2021) are subtle interventions designed to guide users toward making more privacy-conscious decisions without restricting their freedom of choice. They are commonly used in social media platforms, e.g. Facebook’s contextual privacy warnings (Meta, 2024) and Chrome browser-based security warnings (Google, 2025) These nudges typically provide a uniform textual content—such as generic privacy explanations—to all users, regardless of individual differences in privacy literacy. As a result, they risk producing mismatched comprehension: users with limited privacy knowledge may struggle to understand the information (under-comprehension), while more knowledgeable users may find it redundant or oversimplified (over-comprehension), ultimately reducing the effectiveness of the intervention. Hence, they can be personalised based on a user’s expertise or profile, implicitly via Gaze3P.
  • Privacy labels and transparency notices: Similar to privacy nudges, Gaze3P can be used in privacy labels and transparency notices, e.g. Apple’s App Store privacy labels (Apple, 2025), to help users understand and control their privacy choices according to their profiles to avoid mismatched comprehension.
  • Cohort-based recommendations: Cohort-based recommendations, e.g. Google’s Federated Learning of Cohorts (FLoC) (Google, 2021), group users into segments (or cohorts) based on shared characteristics, such as behaviour, preferences, or demographics, and generate recommendations tailored to each group. Instead of personalising for individuals, the system provides suggestions optimised for the typical member of a cohort, balancing personalisation and privacy by avoiding the need for fine-grained individual profiling. Here, again, Gaze3P can be used to profile users according to their gaze behaviour such as privacy preferences.
Appendix D Cognitive Theories and Privacy